OpenStudy (hannahduhe):

if segment YB is parallel to segment ZA find the value of x and the length of segment YB A)x=2/5, YB=3 B)x=2/5,YB=7 C)x=2/3,,YB=3 D)x=2/3,YB=7

OpenStudy (phi):

your picture does not match the question

OpenStudy (phi):

is it clear that the distance between the two parallel lines is the same, no matter where you measure it? in other words, YZ= AB (we could prove the two triangles are congruent and then use corresponding sides of congruent triangles to prove this) start by writing an equation, know YZ = AB

OpenStudy (hannahduhe):

so 16x-4=4-4x

OpenStudy (phi):

yes. to solve add +4x to both sides then add +4 to both sides

OpenStudy (hannahduhe):

so 16x-4+4=4-4x+4

OpenStudy (phi):

ok you add 4 to both sides. now simplify on the left side: -4+4 is 0 on the right side you have 4+4 which is 8 16x -4 +4 = 4 -4x + 4 becomes 16x = -4x + 8 now add+4x to both sides 16x+4x = -4x + 4x + 8 the right side has -4x + 4x (which means you have 4 x's take away 4 x's, so 0 x's) on the left you have 16 x's plus 4 more x's, so 20 x's 20x = 8 to find x, divide both sides by 20 20/20 x = 8/20 or x = 8/20 now use that to find AZ

OpenStudy (phi):

YB is the same size as AZ (YB and AZ are opposite sides of a rectangle, we know they are the same size)

OpenStudy (phi):

btw, x= 8/20 can be simplified by dividing top and bottom by 4 you will get x= 2/5

OpenStudy (hannahduhe):

would it be B